Problem 6-14 At the end of 2020, Shamrock Company is conducting an impairment test and needs to develop a fair value estimate for machinery used in its manufacturing operations. Given the nature of Shamrock’s production process, the equipment is for special use. (No secondhand market values are available.) The equipment will be obsolete in 2 years, and Shamrock’s accountants have developed the following cash flow information for the equipment. Year Net Cash Flow Estimate Probability Assessment 2021 $5,990 40% 8,240 60% 2022 $(450 ) 20% 1,810 60% 3,880 20% Scrap value 2022 $460 50% 850 50% Click here to view factor tables Using expected cash flow and present value techniques, determine the fair value of the machinery at the end of 2020. Use a 6% discount rate. Assume all cash flows occur at the end of the year. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.) Fair value of the machinery at the end of 2020 $---------------
Answer #1

Cash flows in the year 2021 = ($5,990 * 40%) + ($8,240 * 60%) = $2,396 + $4,944 = $7,340

PV of cash flows in the year 2021 = $7,340 * PV at 6% factor for 1 year = $7,340 * 0.9434 = $6,924.56

Cash flows in the year 2022 = (-$450 * 20%) + ($1,810 * 60%) + ($3,880 * 20%) = -$90 + $1,086 + $776 = $1,772

PV of cash flows in the year 2022 = $1,772 * PV at 6% factor for 2 years = $1,772 * 0.8900 = $1,577.08

Scrap Value in the year 2022 = ($460 * 50%) + ($850 * 50%) = $230 + $425 = $655

PV of Scrap Value in the year 2022 = $655 * 0.8900 = $582.95

Total PV of cash flows and scrap value = $6,924.56 + $1,577.08 + $582.95 = $9,084.59

Fair value of the machinery at the end of 2020 = $9,085